OpenStudy (jayden):

Solve for x: (x+4) over 3 =2.

alones (alones):

Like this? \[\frac{ x-4 }{ 3 }=2\]

OpenStudy (jayden):

yes

alones (alones):

Alright so multiply both sides by 3

OpenStudy (jayden):

ok

alones (alones):

\[ \frac{ x-4 }{ 3 }*(3)=(2)*(3)\] \[\ x−4=6\]

alones (alones):

So than just \(\ add~ 4\) to each side \(\ x−4+4=6+4\)

OpenStudy (jayden):

ok

alones (alones):

You would end up with \(\ x=10\)

OpenStudy (jayden):

alright